OpenStudy (anonymous):

Simplify. sqrt 72 x2 sqrt96

OpenStudy (anonymous):

Is this what the equation looks like?\[\sqrt{72}x \sqrt{92}\]

OpenStudy (anonymous):

lol meant to put \[\sqrt{72}(2)\sqrt{92}\]

OpenStudy (anonymous):

No, it looks like \[\sqrt{72} \times2 \sqrt{96}\]

OpenStudy (anonymous):

I have \[48\sqrt{3}....\]

OpenStudy (anonymous):

I have \[48\sqrt{3}....\]

OpenStudy (anonymous):

is that wrong?

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):